OpenStudy (cutestgirlaround):

please hep what's the value of a–2=3+6a/3

OpenStudy (cutestgirlaround):

I think 3

OpenStudy (xxharleyquinnxx):

a=-5 So that means that -5-2=3+6x-5-/3 so your value should be -7. Hope this helps :)

OpenStudy (cutestgirlaround):

so a=-5 and the means a's value is -5 right i'm sorry i'm still trying to understand this stuff

OpenStudy (xxharleyquinnxx):

Are you finding the value of the whole equation or just A?

OpenStudy (phi):

is the problem \[ a-2 = \frac{3+6a}{3} \]?

OpenStudy (cutestgirlaround):

yes

OpenStudy (cutestgirlaround):

just a

OpenStudy (phi):

as you know, the right side can be written \[ \frac{3+6a}{3} = \frac{3}{3} + \frac{6}{3} a \] can you simplify that ?

OpenStudy (cutestgirlaround):

that would 9/3a right

OpenStudy (phi):

what do you get for \[ \frac{3}{3} + \frac{6}{3} a \]?

OpenStudy (phi):

order of operations: first do the divide 3 / 3 is what ? and 6/3 is what ?

OpenStudy (cutestgirlaround):

3/3 is 0 and 6/3 is 2

OpenStudy (cutestgirlaround):

right?

OpenStudy (phi):

6/3 is 2 but anything divided by itself is 1 (not 0)

OpenStudy (cutestgirlaround):

oh sorry I messed up there

OpenStudy (cutestgirlaround):

is 3/3 is 1 and 6/3 is 2?

OpenStudy (phi):

ok so we figured out that \[ \frac{3+6a}{3} = 1+2a \] so now the problem is \[ a-2 = \frac{3+6a}{3} \\ a-2 = 1+2a\] after we simplified the right side

OpenStudy (cutestgirlaround):

ok so what's nect?

OpenStudy (phi):

a-2 = 1+2a I would add -a to both sides. can you do that ?

OpenStudy (cutestgirlaround):

then that would be 2 = 1+ 2 right?

OpenStudy (phi):

you start with a-2 = 1+2a then you write -a on both sides, like this a-a -2 = 1+2a - a

OpenStudy (cutestgirlaround):

ok so then that = 2 = 1+ 2

OpenStudy (phi):

a-a -2 = 1+2a - a a-a is 0 (you have 1 a take away 1 a, leaves no a's) so it is 0 - 2 = 1 + 2a - a 0-2 is just -2 so you have -2 = 1+2 a - a on the right side you have 2 a's take away 1 a. how many a's are left ?

OpenStudy (cutestgirlaround):

none

OpenStudy (cutestgirlaround):

???

OpenStudy (phi):

two apples take away 1 apple leaves 1 apple the same for 2a - a = a so you have -2 = 1+2 a - a becomes -2 = 1 + a finally add -1 to both sides: -2-1 = 1-1 + a

OpenStudy (cutestgirlaround):

so it's -2=1?